(a) State some examples of annuities. (4marks)

(b) Calculate the following present value of annuities:

     (i) RM2,500 a year for 10 years discounted back to the present at 7%. (3 marks)

     (ii) RM70 a year for 3 years discounted back to the present at 3% (3 marks)
Answer #1

a)1)

Examples of Annuities
1.Life Insurance premiums
2.Recurring deposit schemes
3.Monthly pension schemes
4.Mortgage Payments

b)i)

Present Value of Annuity RM 2500 for 10 years @7%
Present Value of RM 1 for 10 years @7%=7.024(As per Annuity Table)
Present Value of Annuity RM 2500=2500*7.024
=2500*7.024
=17560

b)ii)

Present Value of Annuity RM 70 for 3 years @3%
Present Value of RM 1 for 3 years @3%=2.829(As per Annuity Table)
Present Value of Annuity RM 70=70*2.829
=70*2.829
=198.03